OpenStudy (anonymous):

(9/(3x-1)) -(5x/(2x+3)) subtract rational expression

OpenStudy (anonymous):

\[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\]

OpenStudy (anonymous):

what would the denominator look like? I have troubles with the denominator